OpenStudy (anonymous):

is it possible to show using direct method that if n^2 is odd, then n is odd?

OpenStudy (anonymous):

I'm not really a fan of proof by contradiction

OpenStudy (perl):

use contrapositive

OpenStudy (anonymous):

@perl, *direct method* :P

OpenStudy (perl):

i don't know of a direct method, sorry

OpenStudy (anonymous):

direct method is to show P->Q, assume P, then show Q.

OpenStudy (zzr0ck3r):

why are you not a fan of contradiction? Do you use counter examples to show something is not true?

OpenStudy (anonymous):

because i tend to prove directly first before using contradiction

OpenStudy (zzr0ck3r):

sometimes that's not possible

OpenStudy (zzr0ck3r):

many of the theorems you are using to do any of these things were proved in non direct methods, so you are always sort of using a contradiction :)

OpenStudy (zzr0ck3r):

But then again I also will stop the car and ask for directions ;)

Parth (parthkohli):

If \(n^2\) doesn't have 2 as one of its factors, then \(n \) wouldn't either. (Why?)

sammixboo (sammixboo):

I don't know if this says anything that will help, but I saw this section (in image below) in this article http://www.math.uiuc.edu/~hildebr/347.summer14/even-odd-proofs-sol.pdf

OpenStudy (anonymous):

@ParthKohli n^2 = 2k + 1. Ok how does this implie n does not have 2 as a factor?

OpenStudy (zzr0ck3r):

yes but why....

Parth (parthkohli):

It does. If a number doesn't have prime factor \(p_0\) in its prime factors, then why would any of its factors have \(p_0\) in their prime factors?

Parth (parthkohli):

Not sure if this is a direct proof, but it's the best I could come up with...

OpenStudy (zzr0ck3r):

I don't think you can end a proof with a question mark ;P

OpenStudy (zzr0ck3r):

im just playing

OpenStudy (anonymous):

perhaps it's sometimes really impossible to prove directly as @zzr0ck3r said. but I'd still like to think it's always possible though :/

OpenStudy (zzr0ck3r):

I doubt it. I was just saying that some things are.

OpenStudy (anonymous):

\[ n^2 = 2k+1\implies \frac{2k+1}{n}=n \]An odd number divided by an even number is not an integer, therefore \(n\) must be odd.

Parth (parthkohli):

If \(p\) does not divide \(k\), then \(p\) will not divide \(k_0\), a factor of \(k\). Here, p = 2, \(k = n^2\) and \(k_0 = n\).

OpenStudy (anonymous):

h hope that could help you http://zimmer.csufresno.edu/~larryc/proofs/proofs.direct.html

OpenStudy (anonymous):

Another direct proof method would be to examine each case. Either \(n\) is even or odd. If it can't be even, then it must be odd.

OpenStudy (anonymous):

@wio very clever :)

OpenStudy (anonymous):

there is hope that proving directly is always possible. it just take some cleverness :p. Thanks everyone ^^

OpenStudy (perl):

There are many theorems in math that are proven by contradiction. Proving directly is not always possible. For instant, you wouldn't be able to get cantor's transfinite math if you do not use proof by contradiction. Same goes for many analysis results.

OpenStudy (mathmath333):

let n=2m-1,\(m\in \mathbb{Z}\) \(\large\tt \begin{align} \color{black}{n=2m-1\\ n^2=4m^2-4m+1\\ n^2(mod~2)=(4m^2-4m+1)(mod~2)\\ n^2(mod~2)=(4m^2-4m+1)(mod~2)\equiv 1\\ n^2(mod~2)\equiv 1^2\\ n(mod~2)\equiv 1 }\end{align}\) so n is odd

OpenStudy (perl):

how did you go from n^2 (mod 2 ) = 1 to n mod 2 = 1

OpenStudy (mathmath333):

that the theorm if \(a^k \equiv b^k~~ (mod ~n)\) then \(a\equiv b~~ (mod ~n)\)